# Exploring Cross-model Neuronal Correlations in the Context of Predicting Model Performance and Generalizability

### v3

@@ Methods
$S_{i}=\frac{\rho(N_{i},N_{j})}{|\text{layer}(N_{i})-\text{layer}(N_{j})|+1}.$ (3)
The inclusion of the denominator in the formula is crucial, as it accounts for the expected decrease in correlation between neurons that are more distant from each other in the respective architectures. The score is estimated for all neuron pairs. Following that, the average over all $S_{i}$ and $S_{j}$ values determines the correlation score between the two neural networks in question. Given that $\rho$ falls within the interval of -1 to 1, the resultant correlation score similarly spans from -1 to 1. It should be noted that we calculate the correlation function specifically for neurons in the hidden layer, as the last linear layer does not contribute significantly to understanding the network’s behavior and interpretability.
@@
MNIST 0.880 0.856 0.908 0.831 0.416
Fashion-MNIST 0.756 0.822 0.820 0.755 0.317
Kuzushiji-MNIST 0.900 0.769 0.910 0.894 0.519
Afro-MNIST 0.588 0.629 0.730 0.718 0.059
Numerical Results
@@ A. Small neural networks
ResNet-18 1 0.639 0.203 -0.263 0.054
ResNet-34 X 1 0.267 -0.349 0.015
ResNet-50 X X 1 -0.534 -0.407
ResNet-101 X X X 1 -0.688
@@
Table 4 depicts the robustness of FCNs against the black-box attack detailed in Algorithm 1. The impact of the attack on some test images is illustrated in Figure 1. Results for CNN are also provided, showcasing its inferior robustness, given its role in generating the adversaries. In our specific scenario, we selected a norm value of $\epsilon=\frac{10}{255}$ and $N_{ep}=10$. The results indicate that, on average, Net5, Net2, Net4, Net3, and Net1 ranked from vulnerable to robust. Among the FCNs, Net5 exhibited the lowest overall accuracy. Notably, according to Table 2(d), Net5 displayed the weakest correlation with the other networks, aligning with its subpar performance. Conversely, Net3 demonstrated the highest average correlation score with the other networks, potentially attributed to its architectural similarity with the other FCNs.
In Table 3, the correlation between FCNs and CNN is investigated. Here, due to the large number of neurons involved, we had to limit the correlation analysis to only $10$ test data points. Looking at the correlation values between CNN and the different networks, Net3 consistently demonstrates the highest correlation values across multiple datasets. This high correlation between Net3 and CNN suggests that their predictive patterns and feature extraction align closely across various datasets. In addition, it’s noteworthy that Net5 exhibits the lowest correlation among the networks, which aligns with expectations due to its inferior performance and robustness.
@@ B. Large neural networks
The proposed correlation metric was also investigated on large neural networks such as ResNet (He et al. 2016) and DenseNet (Huang et al. 2017). ResNet, short for Residual Network, is a deep neural network architecture that leverages residual blocks and skip connections to enhance training efficiency and performance. The variants of ResNet include ResNet-18, ResNet-34, ResNet-50, ResNet-101, and ResNet-152 (He et al. 2016). DenseNet, short for Dense Convolutional Network, is a deep neural network architecture that leverages dense connections between layers to promote feature reuse and efficient gradient flow, resulting in compact models that excel in various computer vision tasks. The architecture has several variants, namely DenseNet-121, DenseNet-161, DenseNet-169, and DenseNet-201 (Huang et al. 2017). For both cases, we utilized the pretrained weights on ImageNet (Deng et al. 2009) as our baselines. We limited our analysis to just 10 test data points sampled from the validation set of ImageNet due to the time and space complexities inherent in our methodology. Since investigating the correlation on all of the neurons in all layers was not possible, we focused our analysis solely on the initial output of the fourth layer for all ResNets and the primary output of the third transition layer for all DenseNets. A point to be highlighted is that assessing the correlation metric on the final layers often yields superior accuracy compared to other layers, given that these layers encapsulate more profound and meaningful representations. Considering the absolute values of the correlations, the results shown in Table 5 suggest that the most similar networks to ResNet-18, ResNet-34, ResNet-50, ResNet-101, and ResNet-152 are ResNet-34, ResNet-18, ResNet-101, ResNet-152, and ResNet-101 respectively. As demonstrated in Table 6, DenseNet-169, DenseNet-169, DenseNet-201, and DenseNet-169 have the most similarities to DenseNet-121, DenseNet-161, DenseNet-169, and DenseNet-201. This pattern illustrates that networks with roughly the same number of layers often exhibit more profound partial correlations, serving as a validation of the effectiveness of our proposed correlation metric. Furthermore, deeper neural networks exhibit greater robustness against adversarial attacks. Thus, our findings further reveal that a stronger correlation with a robust model correlates with higher robustness.
@@
DenseNet-121 1 0.657 0.664 0.637
DenseNet-161 X 1 0.688 0.631
DenseNet-169 X X 1 0.696
@@ Discussion
  • Architectural size impact: Smaller architectures tend to exhibit higher correlation with other networks due to their simplicity and shared properties with most neural networks. As networks increase in complexity, this correlation diminishes, reflecting the nuanced decision-making processes that diverge from standard neural network behaviors. Calculating partial correlations, especially for the final layers of the networks, can yield more interpretable insights.
  • Layer depth impact: Evaluation of the partial correlation metric on the earlier layers is typically less accurate since the final layers tend to capture more sophisticated and significant representations of the data.

### v4

@@ Methods
$S_{i}=\frac{|\rho(N_{i},N_{j})|}{|\text{layer}(N_{i})-\text{layer}(N_{j})|+1}.$ (3)
The inclusion of the denominator in the formula is crucial, as it accounts for the expected decrease in correlation between neurons that are more distant from each other in the respective architectures. Moreover, it is essential to utilize the absolute value of $\rho$, as a negative correlation can be transformed into a positive correlation when negative weights are applied in the subsequent layer. The score is estimated for all neuron pairs. Following that, the average over all $S_{i}$ and $S_{j}$ values determines the correlation score between the two neural networks in question. Given that $\rho$ falls within the interval of -1 to 1, the resultant correlation score spans from 0 to 1. It should be noted that we calculate the correlation function specifically for neurons in the hidden layer, as the last linear layer does not contribute significantly to understanding the network’s behavior and interpretability.
@@
MNIST 0.921 0.905 0.930 0.893 0.562
Fashion-MNIST 0.412 0.407 0.428 0.453 0.308
Kuzushiji-MNIST 0.925 0.917 0.949 0.942 0.549
Afro-MNIST 0.777 0.794 0.839 0.840 0.348
@@ A. Small neural networks
ResNet-18 1 0.661 0.288 0.312 0.133
ResNet-34 X 1 0.480 0.402 0.258
ResNet-50 X X 1 0.206 0.113
ResNet-101 X X X 1 0.052
@@
Table 4 depicts the robustness of FCNs against the black-box attack detailed in Algorithm 1. The impact of the attack on some test images is illustrated in Figure 1. Results for CNN are also provided, showcasing its inferior robustness, given its role in generating the adversaries. In our specific scenario, we selected a norm value of $\epsilon=\frac{10}{255}$ and $N_{ep}=10$. The results indicate that, on average, Net5, Net2, Net3, Net4, and Net1 ranked from vulnerable to robust. Among the FCNs, Net5 exhibited the lowest overall accuracy. Notably, according to Table 2(d), Net5 displayed the weakest correlation with the other networks, aligning with its subpar performance and larger number of layers. Conversely, Net4 demonstrated the highest average correlation score with the other networks, potentially attributed to its larger number of neurons per layer.
In Table 3, the correlation between FCNs and CNN is investigated. Here, due to the large number of neurons involved, we had to limit the correlation analysis to only $10$ test data points. Looking at the correlation values between CNN and the different networks, Net3 and Net4 consistently demonstrates the highest correlation values across multiple datasets. This high correlation between the networks suggests that their predictive patterns and feature extraction align closely across various datasets. In addition, it’s noteworthy that Net5 exhibits the lowest correlation among the networks, which aligns with expectations due to its inferior performance and robustness.
@@ B. Large neural networks
The proposed correlation metric was also investigated on large neural networks such as ResNet (He et al. 2016) and DenseNet (Huang et al. 2017). ResNet, short for Residual Network, is a deep neural network architecture that leverages residual blocks and skip connections to enhance training efficiency and performance. The variants of ResNet include ResNet-18, ResNet-34, ResNet-50, ResNet-101, and ResNet-152 (He et al. 2016). DenseNet, short for Dense Convolutional Network, is a deep neural network architecture that leverages dense connections between layers to promote feature reuse and efficient gradient flow, resulting in compact models that excel in various computer vision tasks. The architecture has several variants, namely DenseNet-121, DenseNet-161, DenseNet-169, and DenseNet-201 (Huang et al. 2017). For both cases, we utilized the pretrained weights on ImageNet (Deng et al. 2009) as our baselines. We limited our analysis to just 10 test data points sampled from the validation set of ImageNet due to the time and space complexities inherent in our methodology. Since investigating the correlation on all of the neurons in all layers was not possible, we focused our analysis solely on the initial output of the fourth layer for all ResNets and the primary output of the third transition layer for all DenseNets. A point to be highlighted is that assessing the correlation metric on the final layers often yields superior accuracy compared to other layers, given that these layers encapsulate more profound and meaningful representations. The results shown in Table 5 suggest that the most similar networks to ResNet-18, ResNet-34, ResNet-50, ResNet-101, and ResNet-152 are ResNet-34, ResNet-18, ResNet-34, ResNet-34, and ResNet-34 respectively. As demonstrated in Table 6, DenseNet-161, DenseNet-121, DenseNet-201, and DenseNet-121 have the most similarities to DenseNet-121, DenseNet-161, DenseNet-169, and DenseNet-201. This pattern illustrates that networks with roughly the same number of layers often exhibit more profound partial correlations, serving as a validation of the effectiveness of our proposed correlation metric. Furthermore, deeper neural networks exhibit greater robustness against adversarial attacks. Thus, our findings further reveal that a stronger correlation with a robust model correlates with higher robustness.
@@
DenseNet-121 1 0.780 0.718 0.766
DenseNet-161 X 1 0.697 0.748
DenseNet-169 X X 1 0.725
@@ Discussion
  • Architectural size impact: Shallower architectures tend to exhibit higher correlation with other networks due to their simplicity and shared properties with most neural networks. A greater number of neurons per layer can contribute to a higher correlation score, as it allows for the recognition of a broader range of similar patterns with increased neural capacity.
  • Layer depth impact: Calculating partial correlations can yield more interpretable insights. Evaluation of the partial correlation metric on the earlier layers is typically less accurate since the final layers tend to capture more sophisticated and significant representations of the data.